OpenStudy (anonymous):

Given that F(x) = x² + 2, evaluate F(1) + F(5). is the answer 38?

OpenStudy (anonymous):

no, it's 30...

OpenStudy (anonymous):

F(1) = 1² + 2 = 3 F(5) = 5² + 2 = 27 F(1) + F(2)= 30

OpenStudy (anonymous):

plug in 1 into f(x)...then separately plug in 5 into f(x)... add them f(1)= (1)^2 + 2 f(1) = 3 f(5)= 5^2+ 2 f(5) = 27 27+3 = 30